(x+y)* (x+y)*(x+y)= (x+y)^3 laws of exponents
when I use the distributive law...I get x^3 + 3*x^2*y + 3*x*y^2 + y^3 why can't I prove out (x+y)^3? I am missing something

Answer by stanbon(75887)

(x+y)* (x+y)*(x+y)= (x+y)^3 laws of exponents
when I use the distributive law...I get x^3 + 3*x^2*y + 3*x*y^2 + y^3 why can't I prove out (x+y)^3? I am missing something
-------------------------------------------------
Factor x^3 + 3*x^2*y + 3*x*y^2 + y^3
---
Rearrange:
3*x^2*y + 3*x*y^2 + x^3 + y^3
---
Factor to get:
3(xy)(x+y) + (x+y)( x^2-xy+y^2)
(x+y)[3xy + x^2 -xy + y^2]
(x+y)[x^2 +2xy +y^2]
(x+y)(x+y)^2
= (x+y)^3
========================
Cheers,
Stan H.